### v1

@@ Abstract
As Large Language Models (LLMs) are progressively deployed across diverse fields and real-world applications, ensuring the security and robustness of LLMs has become ever more critical. Retrieval-Augmented Generation (RAG) is a cutting-edge approach designed to address the limitations of large language models (LLMs). By retrieving information from the relevant knowledge database, RAG enriches the input to LLMs, enabling them to produce responses that are more accurate and contextually appropriate. It is worth noting that the knowledge database, being sourced from publicly available channels such as Wikipedia, inevitably introduces a new attack surface. RAG poisoning involves injecting malicious texts into the knowledge database, ultimately leading to the generation of the attacker’s target response (also called poisoned response). However, there are currently limited methods available for detecting such poisoning attacks. We aim to bridge the gap in this work. Particularly, we introduce RevPRAG, a flexible and automated detection pipeline that leverages the activations of LLMs for poisoned response detection. Our investigation uncovers distinct patterns in LLMs’ activations when generating correct responses versus poisoned responses. Our results on multiple benchmark datasets and RAG architectures show our approach could achieve 98% true positive rate, while maintaining false positive rates close to 1%. We also evaluate recent backdoor detection methods specifically designed for LLMs and applicable for identifying poisoned responses in RAG. The results demonstrate that our approach significantly surpasses them.
Index Terms:
LLM, RAG Poisoning, Activation.
I Introduction
Recent advances in Large Language Models (LLMs) have remarkably enhanced performance across a diverse set of Natural Language Processing (NLP) tasks, especially in Question-Answering (QA) tasks [1], due to their stunning capabilities to understand and generate texts. Despite these advances, LLMs also face a range of issues arising from knowledge limitations. On one hand, LLMs are trained on historical data and lack up-to-date knowledge (e.g., GPT-4’s data cutoff is December 2023 [2], which limits its capability to effectively process information beyond that point). On the other hand, they also exhibit knowledge gaps in specialized domains, such as medicine, finance and privacy-related contexts. Due to the absence of relevant knowledge, LLMs may produce inaccurate content and exhibit “hallucination” behaviors, where they generate information that is incorrect, misleading, or fabricated, yet appears plausible and authoritative. These limitations pose challenges to LLM applications in fields such as healthcare [3], economics [4], and scientific research [5, 6].
@@ I Introduction
Retrieval-Augmented Generation (RAG) [7] has emerged as an effective solution that leverages retrievers to incorporate external databases, enriching the knowledge of LLMs and ultimately enabling the generation of up-to-date and accurate responses. RAG comprises three components: knowledge database, retriever, and LLM. The knowledge database consists of a large amount of texts collected from sources such as latest Wikipedia entries [8], new articles [9] and financial documents [4]. The retriever is primarily responsible for using a text encoder (e.g., BERT [10]) to compute the embedding vectors of the user’s question (e.g., “What is the name of the highest mountain?”) and the texts in the knowledge database and selects the top-k retrieved texts from the knowledge database based on similarity. The LLM generates the final response (e.g., “Everest”) based on the user’s question and the retrieved texts. Due to RAG’s powerful knowledge integration capabilities, it has demonstrated impressive performance across a range of QA-like knowledge-intensive tasks [11, 12]. Fig. 1 visualizes an example of RAG.
@@
RAG poisoning refers to the act of injecting malicious or misleading content into the knowledge database, contaminating the retrieved texts in RAG and ultimately leading it to produce the attacker’s desired response (e.g., the target answer could be “Fuji” when the target question is “What is the name of the highest mountain?”). This attack leverages the dependency between LLMs and the knowledge database, transforming the database into a new attack surface to facilitate poisoning. PoisonedRAG  [13] demonstrates the feasibility of RAG poisoning by injecting a small amount of maliciously crafted texts into the knowledge database utilized by RAG. This manipulation effectively compels the model to generate the attacker’s intended response, referred to as a “poisoned response”. Furthermore, injecting a substantial amount of adversarial texts into the knowledge database can mislead the retriever into prioritizing such content, causing it to generalize poisoned responses to queries from unseen domains [14].
The rise of such attacks has drawn significant attention to the necessity of designing robust and resilient RAG systems. For example, SELF-RAG [15] introduces the use of LLM self-reflection to evaluate the usefulness of the retrieved texts; INSTRUCTRAG [16] utilizes LLMs to analyze how to extract correct answers from noisy retrieved documents; RobustRAG [17] introduces multiple LLMs to generate answers from the retrieved texts, and then aggregates the responses. However, the aforementioned defense methods necessitate the integration of additional large models, incurring considerable overheads. Moreover, they’re all “best-effort” schemes, offering no guarantee on the defense effectiveness. Once an attacker becomes aware of the defense strategy, they can launch attacks that remain imperceptible to the schemes. Consequently, the user cannot reliably ascertain whether the response generated by such schemes is trustworthy.
@@
In our work, we shift our focus to leverage the intrinsic properties of LLMs for detecting RAG poisoning, rather than relying on external models. Our view is that if we can accurately determine whether a RAG’s response is correct or poisoned, we can effectively thwart RAG poisoning attacks. We attempt to observe LLM’s answer generation process to determine whether the response is compromised or not. Our focus is not on detecting malicious inputs to LLMs, as we consider the consequences of malicious responses to be far more detrimental and indicative of an attack. The growing body of research on using activations to explain and control LLM behavior [18, 19] has inspired us to leverage LLMs’ activations to determine RAG’s responses are poisoned answers or not. Specifically, we empirically analyze the activations of the final token in the input sequence across all layers of the LLM. Our findings demonstrate that it is highly feasible to differentiate between these two types of responses by comparing the activations of the LLM when generating correct responses versus poisoned ones. Based on this, we propose a systematic and automated detection pipeline, namely RevPRAG, which consists of three key components: poisoned data collection, activation collection, and the design of detection model, as illustrated in Fig. 2. It is important to note that this detection method will not alter the RAG workflow or weaken its performance, thereby offering superior adversarial robustness compared to methods that rely solely on filtering retrieved texts.
To evaluate our approach, we systematically demonstrate the effectiveness of RevPRAG across various LLM architectures, including GPT2-XL-1.5B, Llama2-7B, Mistral-7B, Llama3-8B, and Llama2-13B, as well as diverse retrievers such as Contriever, Contriever-ms, DPR-mul, and ANCE. RevPRAG performs consistently well, achieving over 98% accuracy across three different datasets. We also conduct extensive ablation studies to investigate the impact of different components and parameters on the RevPRAG’s performance. It is widely recognized that LLMs can generate non-factual responses due to hallucinations; however, such responses are fundamentally different from the poisoned responses resulting from poisoning attacks. To further differentiate whether an incorrect response from RAG is poisoned or non-factual, we also construct corresponding datasets for training and testing. Experimental results show that our proposed method can clearly distinguish between poisoned and non-factual responses as well.
In summary, our contributions are as follows:
@@
    We introduce RevPRAG, a novel and automated pipeline for detecting whether a RAG’s response is a poisoned or not. By supporting the construction of new datasets to reflect emerging RAG attacks, RevPRAG is capable of detecting new RAG threats.
@@
II Background and Related Work
II-A Retrieval Augmented Generation
The LLM community has been challenged by the generation of inaccurate and outdated responses due to their incomplete and static knowledge, and RAG [20, 7] was introduced as a promising ameliorative paradigm to provide more accurate, relevant, and up-to-date external information. RAG is most striking when it leverages the fact that the knowledge database can be seamlessly refreshed with newly added content. This feature allows the system to efficiently keep up with dynamic knowledge domains, eliminating the need for costly and time-consuming LLM fine-tuning. Because of its great flexibility, RAG arises in innumerable industrial contexts (e.g., ChatGPT Retrieval Plugin [21], WikiChat [22], FinGPT [23], and ChatRTX [24]) as well as in specialized domains (e.g., medical diagnostic assistants [25] and email/code completion [26]).
RAG comprises three components: knowledge database, retriever, and LLM. The knowledge database contains data that is either newly added or updated the existing information in the LLM’s training set, often collected from sources such as Wikipedia [8] and other platforms. We define the knowledge database as $\mathcal{C}=\{p_{1},p_{2},\cdots,p_{m}\}$ where $p_{i}$ represents the $i$-th piece of collected information. The retriever identifies the $k$ texts most similar to the query $q$ from the knowledge database based on similarity comparison, providing them as external information. The LLM then generates an answer by simultaneously processing the query $q$ and the top-$k$ texts.
The workflow of RAG, having introduced its components, deserves further explanation. As illustrated in Fig. 1, it consists of two main steps: retrieval step and generation step.
In the retrieval step, the retriever generates the top $k$ most relevant pieces of knowledge for the query $q$. First, we employ two encoders, $E_{q}$ and $E_{p}$, which can either be identical or radically different. Encoder $E_{q}$ is responsible for transforming the user’s query $q$ into an embedding vector $E_{q}(q)$, while encoder $E_{p}$ is designed to convert all the information $p_{i}$ in the knowledge database into embedding vectors $E_{p}(p_{i})$. For each $E_{p}(p_{i})$, the similarity with the query $E_{q}(q)$ is computed using $\text{sim}(E_{q}(q),E_{p}(p_{i}))$, where $\text{sim}(\cdot,\cdot)$ quantifies the similarity between two embedding vectors, such as cosine similarity or the dot product. Finally, the top $k$ most relevant pieces are selected as the external knowledge $\mathcal{C}_{q}$ for the query $q$.
The generation step is dedicated to generating a response $\text{LLM}(q,\mathcal{C}_{q})$ based on the query $q$ and the relevant information $\mathcal{C}_{q}$. First, we combine the query $q$ and the external knowledge $\mathcal{C}_{q}$ using a standard prompt (see Fig. 5 for the complete prompt). Taking advantage of such a prompt, the LLM generates an answer $\text{LLM}(q,\mathcal{C}_{q})$ to the query $q$ and can quickly reference the information in the knowledge database. Therefore, RAG is a significant accomplishment, as it addresses the limitations of LLMs in acquiring up-to-date and domain-specific information.
II-B Inner States Analysis of LLM
The interpretability of LLMs has been explored in a collection of works inspired by their inner states and latent spaces. GemmaScope [27] leverages sparse autoencoders to analyze the latent spaces of LLMs, revealing meaningful features within the model activations. These latent states are also capable of encoding safety concerns [28] and task drift [29]. BEEAR [30] detects backdoors by identifying uniform embedding drifts triggered by specific malicious training-time triggers. These methods facilitate the construction of probes to predict the behavior of LLMs, such as hallucinations [31, 32] and task drift [29], as well as to reveal latent knowledge [33].
The logit lens leverages the model’s unembedding matrix to decode representations from intermediate layers, providing valuable insights into its decision-making process and facilitating the mitigation of harmful behaviors. The decoding outputs of intermediate layers are analyzed to identify human-understandable concepts encoded in the parameter vectors of the transformer’s feed-forward layers, enabling the manipulation of model outputs to mitigate harmful content [34]. Late-layer MLPs frequently reduce the probability of the maximum-likelihood token, and this reduction can be self-corrected by removing certain attention layers [35]. Intermediate layers are decoded to reveal how harmful outputs are mimicked by language models, identifying decision-making shifts and suggesting the removal of key layers to enhance output accuracy against misleading inputs [36]. Furthermore, LLM Factoscope [19] employs the logit lens to decode intermediate layer outputs, capturing changes in output indices and probabilities within the LLM.
II-C Retrieval Corruption Attack
Due to the growing attention on RAG, attacks on RAG have also been widely studied. RAG can improperly generate answers that are severely impacted or compromised once the knowledge database is contaminated [13, 37, 38]. Furthermore, the retriever can also be vulnerable to attacks, which may result in the selection of irrelevant or misleading external information, without even modifying the knowledge database [39, 40]. Specifically, an attacker can inject a small amount of malicious information onto a website, which is then retrieved by RAG [41]. PoisonedRAG [13] injects malicious text into the knowledge database and formalizes the knowledge poisoning attack as an optimization problem, thereby enabling the LLM to generate target responses selected by the attacker. GARAG [42] was introduced to provide low-level perturbations to RAG and to search for adversarial text targeting both components of RAG simultaneously. PRCAP [14] injects adversarial samples into the knowledge database, where these samples are generated by perturbing discrete tokens to enhance their similarity with a set of training queries. These methods have yielded striking attack results, and in our work, we have selected several state-of-the-art attack methods as our base attacks on RAG.
II-D The Robustness of RAG
Efforts have been made to develop defenses in response to poisoning attacks and noise-induced disruptions, and significant accomplishments have been achieved. As a result, the robustness of RAG, defined by its capacity to resist poisoning attacks and integrate accurate external information, plays a prominent role in practical applications. Adversarial texts are applied to poison the knowledge database [14, 13, 39], necessitating rigorous scrutiny before feeding the retrieved information into the LLM to prevent such adversarial attacks and out-of-distribution data [43, 44, 45, 46]. RobustRAG [17] mitigates the impact of poisoned texts through a voting mechanism, while INSTRUCTRAG [16] explicitly learns the denoising process to address poisoned and irrelevant information. Other approaches to enhance robustness include prompt design [47, 48], plug-in models [49], and specialized models [50, 15]. While these methods alleviate the issues of poisoning and noise, they share similar limitations—specifically, that adaptive attackers can often circumvent the defenses once the underlying algorithms are made publicly available [51, 52, 53, 54, 55]. In this work, we present a detection perspective on defending against RAG attacks, which allows for the open sharing of detection algorithms and models without the risk of being easily circumvented by adaptive attacks.
III PRELIMINARY
In this section, we first define our threat model, and present our design objectives. We then detail the rationale behind detecting poisoning attacks through activation analysis.
III-A Threat Model
Attacker’s Goal. Assume that the attacker preselects a target question set $Q$, consisting of $q_{1},q_{2},\cdots,q_{n}$, and the corresponding target answer set $A$, represented as $a_{1},a_{2},\cdots,a_{n}$. The attacker’s goal is to compromise the RAG system by contaminating the retrieval texts, thereby manipulating the LLM to generate the target response $a_{i}$ for each query $q_{i}$. Taking the upper half of Fig. 2 as an example, the attacker’s target question $q_{i}$ is “What is the name of the highest mountain?”, with the target answer being “Fuji”. To achieve this, the attacker injects malicious texts into the knowledge database $C$, effectively contaminating its content as a source of information for RAG. Consequently, the LLM generates the incorrect target answer “Fuji” based on the retrieved poisoned texts.
Attacker’s Capabilities. We assume that an attacker can inject $m$ poisoned texts $P$ for each target question $q_{i}$, represented as $p_{i}^{1},p_{i}^{2},...,p_{i}^{m}$. The attacker does not possess knowledge of the LLM utilized by the RAG, but has white-box access to the RAG retriever. This assumption is reasonable, as many retrievers are openly accessible on platforms like HuggingFace, such as Contriever [56], Contriever-ms (fine-tuned on MS-MARCO) [56], and ANCE [57]. The poisoned texts can be integrated into the RAG’s knowledge database through two ways: the attacker publishing the maliciously crafted content on open platforms like Wikipedia, or utilizing data collection agencies to disseminate the poisoned texts.
III-B Design Objectives
Accuracy. We aim to develop an accurate model for detecting poisoned responses. Accuracy, in this context, requires identifying all poisoned responses while keeping false rates to a minimum. A high false positive rate would incorrectly flag many valid responses as poisoned, disrupting the model’s functionality and reducing its effectiveness. Conversely, a high false negative rate would allow poisoned responses to go undetected. This work focuses on minimizing both false positives and false negatives to ensure reliable detection.
Non-intrusiveness. Another key objective is to preserve the integrity of the RAG system by avoiding modifications to its core architecture. Our aim is to design a detection model that is both highly accurate and easy to integrate. To ensure practicality, we steer clear of resource-intensive changes, such as fine-tuning the large model, which can be costly. Instead, we focus on achieving accurate detection of poisoned responses within the existing RAG workflow, leveraging available information without introducing additional overhead.
III-C Rationale
The rationale behind our research stems from the conflict between the poisoned texts in the knowledge database and the knowledge originally learned by the LLMs. This conflict arises as attackers seek to manipulate the models into generating poisoned responses that contradict accurate information. This inherent conflict drives our efforts to develop methods for detecting and exploiting it. The activations of LLMs represent input data at varying layers of abstraction, enabling the model to progressively extract high-level semantic information from low-level features. The extensive information encapsulated in these activations comprehensively reflects the entire decision-making process of the LLM. Consequently, we posit that activations can, to a certain extent, capture the aforementioned conflict. Fig. 3 illustrates the t-SNE [58] representation of the mean activations across all layers for Mistral-7B and Llama2-7B on the NQ dataset, comparing poisoned and correct responses. The results clearly highlight the distinguishability between correct and poisoned responses. In summary, activations from all layers of the LLM can effectively serve as indicators for detecting poisoned responses in RAG systems.
IV Methodology
This section begins by providing an overview of our pipeline. We then describe the methodology for crafting effective poisoned texts to execute attacks. Following the successful influence of the poisoned texts on the LLM’s responses, we collect and process the corresponding activations. Finally, we present the architectural design of the RevPRAG model.
IV-A Overview of the method
We introduce RevPRAG, a pipeline designed to leverage LLM activations for detecting knowledge poisoning attacks in RAG systems, as illustrated in Fig. 2. This pipeline enables seamless expansion of dataset coverage for the poisoning attack detection model.
To detect poisoning attacks against RAG, we begin by collecting the texts used in the attack. These poisoned texts are injected into the RAG system’s knowledge database, contaminating the previously clean retrieval texts. Using the poisoned texts as a basis, we label the activation data based on whether the LLM generates the attacker’s target response. Additionally, we capture the LLM’s activations for all queries.
After collecting and preprocessing the data, we train a Siamese network-based model. This approach minimizes the embedding distance for data within the same class while maximizing the distance between data from different classes, ensuring robust detection of poisoning attacks.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) enriches the input to LLMs by retrieving information from the relevant knowledge database, enabling them to produce responses that are more accurate and contextually appropriate. It is worth noting that the knowledge database, being sourced from publicly available channels such as Wikipedia, inevitably introduces a new attack surface. RAG poisoning attack involves injecting malicious texts into the knowledge database, ultimately leading to the generation of the attacker’s target response (also called poisoned response). However, there are currently limited methods available for detecting such poisoning attacks. We aim to bridge the gap in this work by introducing RevPRAG, a flexible and automated detection pipeline that leverages the activations of LLMs for poisoned response detection. Our investigation uncovers distinct patterns in LLMs’ activations when generating poisoned responses versus correct responses. Our results on multiple benchmarks and RAG architectures show our approach can achieve a 98% true positive rate, while maintaining a false positive rate close to 1%.
RevPRAG: Revealing Poisoning Attacks in Retrieval-Augmented Generation through LLM Activation Analysis
Xue Tan1,2, Hao Luan1,2, Mingyu Luo1,2, Xiaoyan Sun2, Ping Chen1, Jun Dai2 School of Computer Science, Fudan University, Shanghai, China1 {xuetan23, hluan23}@m.fudan.edu.cn, luomingyu2002@gmail.com Department of Computer Science, Worcester Polytechnic Institute, MA, USA2
1 Introduction
@@
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) has emerged as an effective solution that leverages retrievers to incorporate external databases, enriching the knowledge of LLMs and ultimately enabling the generation of up-to-date and accurate responses. RAG comprises three components: knowledge database, retriever, and LLM. Fig. 1 visualizes an example of RAG. The knowledge database consists of a large amount of texts collected from sources such as latest Wikipedia entries Thakur et al. (2021), new articles Soboroff et al. (2018) and financial documents Loukas et al. (2023). The retriever is primarily responsible for retrieving the texts that are most related to the user’s query from the knowledge database. These texts will later be fed to LLM as a part of the prompt to generate responses (e.g., “Everest") for users’ queries (e.g., “What is the name of the highest mountain?"). Due to RAG’s powerful knowledge integration capabilities, it has demonstrated impressive performance across a range of QA-like knowledge-intensive tasks Lazaridou et al. (2022); Jeong et al. (2024).
@@
RAG poisoning refers to the act of injecting malicious or misleading content into the knowledge database, contaminating the retrieved texts in RAG and ultimately leading it to produce the attacker’s desired response (e.g., the target answer could be “Fuji" when the target question is “What is the name of the highest mountain?"). This attack leverages the dependency between LLMs and the knowledge database, transforming the database into a new attack surface to facilitate poisoning. PoisonedRAG  Zou et al. (2024) demonstrates the feasibility of RAG poisoning by injecting a small amount of maliciously crafted texts into the knowledge database utilized by RAG. The rise of such attacks has drawn significant attention to the necessity of designing robust and resilient RAG systems. For example, INSTRUCTRAG Wei et al. (2024) utilizes LLMs to analyze how to extract correct answers from noisy retrieved documents; RobustRAG Xiang et al. (2024) introduces multiple LLMs to generate answers from the retrieved texts, and then aggregates the responses. However, the aforementioned defense methods necessitate the integration of additional large models, incurring considerable overheads. Meanwhile, it is difficult to promptly assess whether the current response of RAG is trustworthy or not.
@@
In our work, we shift our focus to leverage the intrinsic properties of LLMs for detecting RAG poisoning, rather than relying on external models. Our view is that if we can accurately determine whether a RAG’s response is correct or poisoned, we can effectively thwart RAG poisoning attacks. We attempt to observe LLM’s answer generation process to determine whether the response is compromised or not. It is worth noting that our focus is not on detecting malicious inputs to LLMs, as we consider the consequences of malicious responses to be far more detrimental and indicative of an attack. The growing body of research on using activations to explain and control LLM behavior Ferrando et al. (2024); He et al. (2024) provides us inspiration. Specifically, we empirically analyze the activations of the final token in the input sequence across all layers of the LLM. Our findings demonstrate that it is highly feasible to differentiate between these two types of responses by comparing the activations of the LLM when generating correct responses versus poisoned ones. Based on this, we propose a systematic and automated detection pipeline, namely RevPRAG, which consists of three key components: poisoned data collection, LLM activation collection and preprocessing, and the detection model design. It is important to note that this detection method will not alter the RAG workflow or weaken its performance, thereby offering superior adversarial robustness compared to methods that rely solely on filtering retrieved texts.
To evaluate our approach, we systematically demonstrate the effectiveness of RevPRAG across various LLM architectures, including GPT2-XL-1.5B, Llama2-7B, Mistral-7B, Llama3-8B, and Llama2-13B. RevPRAG performs consistently well, achieving over 98% true positive rate across different datasets.
Our contributions can be summarized as follows:
@@
    We introduce RevPRAG, a novel and automated pipeline for detecting whether a RAG’s response is poisoned or not. To address emerging RAG poisoning attacks, RevPRAG allows new datasets to be constructed accordingly for training the model, enabling effective detection of new threats.
@@
2 Background and Related Work
2.1 Retrieval Augmented Generation
RAG comprises three components: knowledge database, retriever, and LLM. As illustrated in Fig. 1, RAG consists of two main steps: retrieval step and generation step. In the retrieval step, the retriever acquires the top $k$ most relevant pieces of knowledge for the query $q$. First, we employ two encoders, $E_{q}$ and $E_{p}$, which can either be identical or radically different. Encoder $E_{q}$ is responsible for transforming the user’s query $q$ into an embedding vector $E_{q}(q)$, while encoder $E_{p}$ is designed to convert all the information $p_{i}$ in the knowledge database into embedding vectors $E_{p}(p_{i})$. For each $E_{p}(p_{i})$, the similarity with the query $E_{q}(q)$ is computed using $\text{sim}(E_{q}(q),E_{p}(p_{i}))$, where $\text{sim}(\cdot,\cdot)$ quantifies the similarity between two embedding vectors, such as cosine similarity or the dot product. Finally, the top $k$ most relevant pieces are selected as the external knowledge $\mathcal{C}_{q}$ for the query $q$. The generation step is to generating a response $\text{LLM}(q,\mathcal{C}_{q})$ based on the query $q$ and the relevant information $\mathcal{C}_{q}$. First, we combine the query $q$ and the external knowledge $\mathcal{C}_{q}$ using a standard prompt (see Fig. 5 for the complete prompt). Taking advantage of such a prompt, the LLM generates an answer $\text{LLM}(q,\mathcal{C}_{q})$ to the query $q$. Therefore, RAG is a significant accomplishment, as it addresses the limitations of LLMs in acquiring up-to-date and domain-specific information.
2.2 Retrieval Corruption Attack
Due to the growing attention on RAG, attacks on RAG have also been widely studied. RAG can improperly generate answers that are severely impacted or compromised once the knowledge database is contaminated Zou et al. (2024); Xue et al. (2024); Jiao et al. (2024). Specifically, an attacker can inject a small amount of malicious information onto a website, which is then retrieved by RAG Greshake et al. (2023). PoisonedRAG Zou et al. (2024) injects malicious text into the knowledge database, and formalizes the knowledge poisoning attack as an optimization problem, thereby enabling the LLM to generate target responses selected by the attacker. GARAG Cho et al. (2024) was introduced to provide low-level perturbations to RAG. PRCAP Zhong et al. (2023) injects adversarial samples into the knowledge database, where these samples are generated by perturbing discrete tokens to enhance their similarity with a set of training queries. These methods have yielded striking attack results, and in our work, we have selected several state-of-the-art attack methods as our base attacks on RAG.
2.3 The Robustness of RAG
Efforts have been made to develop defenses in response to poisoning attacks and noise-induced disruptions. RobustRAG Xiang et al. (2024) mitigates the impact of poisoned texts through a voting mechanism, while INSTRUCTRAG Wei et al. (2024) explicitly learns the denoising process to address poisoned and irrelevant information. Other approaches to enhance robustness include prompt design Cho et al. (2023); Press et al. (2023), plug-in models Baek et al. (2023), and specialized models Yoran et al. (2023); Asai et al. (2023). However, these methods may, on one hand, rely on additional LLMs, leading to significant overhead. On the other hand, they primarily focus on defense mechanisms before the LLM generates a response, making it challenging for these existing approaches to detect poisoning attacks in real-time while the LLM is generating the response Athalye et al. (2018); Bryniarski et al. (2021); Carlini and Wagner (2017); Carlini (2023); Tramer et al. (2020). In this work, to defend against RAG attacks, we present a detection mechanism that can promptly capture key information during the model’s response generation process and assess whether the current response is trustworthy or not.
3 PRELIMINARY
3.1 Threat Model
Attacker’s goal. We assume that the attacker preselects a target question set $Q$, consisting of $q_{1},q_{2},\cdots,q_{n}$, and the corresponding target answer set $A$, represented as $a_{1},a_{2},\cdots,a_{n}$. The attacker’s goal is to compromise the RAG system by contaminating the retrieval texts, thereby manipulating the LLM to generate the target response $a_{i}$ for each query $q_{i}$. For example, the attacker’s target question $q_{i}$ is “What is the name of the highest mountain?", with the target answer being “Fuji".
Attacker’s capabilities. We assume that an attacker can inject $m$ poisoned texts $P$ for each target question $q_{i}$, represented as $p_{i}^{1},p_{i}^{2},...,p_{i}^{m}$. The attacker does not possess knowledge of the LLM utilized by the RAG, but has white-box access to the RAG retriever. This assumption is reasonable, as many retrievers are openly accessible on platforms like HuggingFace. The poisoned texts can be integrated into the RAG’s knowledge database through two ways: the attacker publishing the malicious content on open platforms like Wikipedia, or utilizing data collection agencies to disseminate the poisoned texts.
@@ 3.1 Threat Model
3.2 Rationale
The activations of LLMs represent input data at varying layers of abstraction, enabling the model to progressively extract high-level semantic information from low-level features. The extensive information encapsulated in these activations comprehensively reflects the entire decision-making process of the LLM. The activations has been applied to factual verification of the output content  He et al. (2024) and detection of task drift Abdelnabi et al. (2024). Due to the fact that LLM produces different activations when generating varying responses, we hypothesize that LLM will also exhibit distinct activations when generating poisoned responses compared to correct ones. Fig. 2 presents the visualizations of activations for correct and poisoned responses using t-SNE (t-Distributed Stochastic Neighbor Embedding). It visualizes the mean activations across all layers for two LLMs, Mistral-7B and Llama2-7B, on the Natural Questions dataset. This clearly demonstrates the distinguishability between the two types of responses, to some extent, supports our conjecture.